OpenStudy (anonymous):

Which of the following is an equivalent form of the compound inequality −22 > −5x − 7 greater than or equal to −3? −5x − 7 < −22 and −5x − 7 greater than or equal to −3 −5x − 7 > −22 and −5x − 7 greater than or equal to −3 −5x > −22 and −7 greater than or equal to −3 −5x − 7 < −22 and −5x − 7 less than or greater to −3

jimthompson5910 (jim_thompson5910):

−22 > −5x − 7 >= −3 breaks down into −22 > −5x − 7 and -5x - 7 >= -3 and you can flip the first inequality to get -5x - 7 < -22 and -5x - 7 >= -3

OpenStudy (anonymous):

THANK YOU!!! :)

jimthompson5910 (jim_thompson5910):

you're welcome